# FARM: Few-shot Adaptive Malware Family Classification under Concept Drift

### v1

@@ V-C Comparison with Baselines
To contextualize FARM’s performance, we compare it against CADE [47], as well as two ablated variants of our approach. The first, Percentile, uses a triplet autoencoder with class-wise mean centroids and distance thresholds defined by the 95th percentile. The second, FARM, is our proposed full approach, which combines metric learning with DBSCAN-based class modeling and dynamic thresholding per cluster. All models are trained using the same encoder architecture and training settings described in  Section IV-B, ensuring a consistent basis for comparison.
@@
Comparative results are presented in Table V. Our approach, FARM, outperforms the state-of-the-art method as well as the ablated Percentile variant, achieving a stronger balance between precision and recall. The Percentile model uses a triplet autoencoder with class-wise centroids and fixed distance thresholds based on the 95th percentile, while FARM introduces DBSCAN-based clustering and dynamic thresholding to more accurately represent intra-family variations. These results demonstrate that incorporating unsupervised clustering improves class boundary modeling and enhances performance in the presence of label drift.The reported values in the table represent the mean and standard deviation of precision, recall, and F1 score, measured across experiments on all eight unseen malware families. In each evaluation run, one novel family is introduced at a time, and performance is assessed independently.
@@ VI-A Experimental Setup
Once such a cluster is formed, its centroid is treated as a class prototype and incorporated into the classification model. In the case of label drift, where the cluster represents a previously unseen malware family, this prototype is treated as a novel class; allowing the system to classify new instances with only a few examples. In contrast, for covariate drift, where the cluster corresponds to behavioral or structural variations within a previously known malware family, the new prototype is linked to the corresponding existing family label in the training set. In other words, instead of creating a new class, the prototype is added as an additional representative of that known malware family, enriching its characterization in the embedding space. In both scenarios, classification is performed using nearest-prototype matching in the latent space.
@@ VI-C Adaptation under Label Drift
To evaluate the system’s ability to adapt to label drift, we simulate the emergence of a novel malware families that were entirely excluded from training phase. Samples from each unseen family are introduced one at a time. When a sample is identified as drifted, based on its distance from all known class prototypes, it is stored in a buffer designated for potential novel class discovery. Once the buffer accumulates enough drifted samples, DBSCAN is applied in the latent space to identify cohesive clusters. If a valid cluster is detected, its centroid is promoted to a new class prototype and incorporated into the classifier. This enables the system to recognize future samples from that family without requiring full model retraining.
@@
VI-C1 Generalization accuracy on unseen classes:
To evaluate the generalization capability of the model to previously unseen classes, we conduct standard few-shot classification experiments under multiple N-way K-shot settings. This setup is commonly used in the few-shot learning literature to simulate realistic scenarios where where a model must learn to distinguish among N new classes using only K labeled examples per class [40, 6, 42].
In our case, each N-way episode includes N unseen malware families, randomly selected from the eight families excluded during training. For each family, K labeled samples are provided as support, and 15 additional samples are drawn to form the query set. This simulates a deployment scenario where the model encounters a small number of labeled examples from emerging threats and must quickly generalize to classify related variants.We follow the experimental setup in [40], running 600 episodes for each N-way K-shot configuration. The mean classification accuracy and 95% confidence intervals reported in Table VIII are computed over these 600 episodes. This episodic evaluation captures the variability in task difficulty and provides a statistically grounded estimate of the model’s generalization performance.
As shown in Table VIII, FARM achieves strong performance across all configurations. Accuracy improves as the number of support samples per class (K) increases, although gains diminish beyond the 10-shot setting. Conversely, increasing the number of classes (N) makes the task more challenging, resulting in lower accuracy due to finer-grained decision boundaries required among a greater variety of families.
These results address RQ3, showing that FARM maintains strong generalization across various few-shot classification settings.
@@ D. Retraining for Label Drift Integration
While few-shot adaptation enables fast and lightweight incorporation of novel malware families, we also implement a retraining strategy to enhance long-term model integration. This approach is used exclusively in the context of label drift, as the number of drifted samples detected under covariate drift is typically too low to support stable retraining. To avoid overfitting and catastrophic forgetting, retraining is only triggered once the drift buffer accumulates at least 100 samples from a novel class. This threshold is defined relative to the scale of the original training data, balancing the need to initiate retraining early enough to remain responsive to meaningful distribution shifts, while ensuring that enough samples are available to support stable gradient updates. Although this value is empirically chosen, it reflects a trade-off that can be adjusted based on the model’s complexity, the volume of original training data, or observed drift dynamics. Once the threshold is met, the model is retrained using both the original training data and the buffered samples, allowing the encoder to update its latent space to reflect the new distribution. After retraining, DBSCAN is reapplied to establish new clusters and thresholds, maintaining the same classification and drift detection framework.
Table IX compares the performance of the prototype-based few-shot adaptation and the retrained model across multiple unseen malware families. The retrained model consistently improves generalization, boosting the average F1 score from 0.85 to 0.94, a relative improvement of 10.6%. This highlights the complementary nature of few-shot adaptation for rapid response and full retraining for deeper integration into the classification system.
TABLE IX: Comparison of prototype-based adaptation and full retraining on unseen malware families (label drift scenario).
@@ VII Discussion
FARM’s performance demonstrates the strength of combining triplet-based metric learning with unsupervised clustering and prototype-based adaptation to handle both covariate and label drift in malware classification. A central component of the system is DBSCAN clustering in the latent space, which enables the discovery of coherent malware subgroups without requiring predefined family class counts. However, the clustering performance is sensitive to the selection of hyperparameters such as $\varepsilon$ and minPts. In this work, $\varepsilon$ was determined using a k-distance elbow heuristic [38], and minPts was set proportional to the latent dimension. While this provides a practical and automated thresholding method, more adaptive or data-driven parameter tuning may lead to improved drift detection sensitivity, especially in complex or noisy regions of the feature space.
An illustrative case of this sensitivity is the vobfus family, which exhibited lower drift detection accuracy due to behavioral overlap with training-set families like fareit and zbot. This overlap caused vobfus samples to be embedded near existing families, resulting in frequent misclassifications. Such cases highlight a limitation in the current embedding space’s ability to fully disentangle semantically similar but distinct behaviors. Another factor influencing representation quality is the construction of training triplets. In the current approach, anchor-positive-negative combinations are sampled randomly without employing strategies like hard positive or hard negative mining. While this simplifies training and prevents overfitting to extreme cases, it may limit the model’s ability to fine-tune decision boundaries in ambiguous or tightly clustered regions.
Overall, the design of FARM effectively balances scalability and adaptability in dynamic malware environments. Yet, these observations suggest that improvements in both clustering calibration and triplet selection strategies could further enhance the system’s robustness, especially in edge-case drift scenarios or where latent overlaps challenge separability.
@@ VIII Conclusion
This work addresses a critical challenge in malware classification, maintaining model effectiveness amid continuous concept drift. By integrating metric learning, few-shot adaptation, and conditional retraining, FARM provides a principled and adaptive framework for identifying and responding to both evolved and novel malware families. Our use of a triplet autoencoder enables reliable drift detection, which surpasses the state of the art, without reliance on softmax confidence scores, while prototypical networks support fast, low-cost adaptation using minimal labeled data. Furthermore, the inclusion of a buffer-triggered retraining mechanism allows FARM to consolidate new knowledge into the model, refreshing its latent representation and enhancing long-term performance. Experimental results confirm that FARM improves classification across both covariate and label drift scenarios, demonstrating strong generalization and scalability. These qualities position FARM as a practical and forward-looking solution for real-world malware detection systems operating in adversarial and temporally evolving environments. Extensions to this work may include handling packed malware, which introduces additional drift through obfuscation and presents challenges for static analysis. Similarly, enhancing the interpretability of drift decisions by identifying the underlying changes that trigger detection could improve the system’s transparency and support more informed responses in operational settings.

### v2

@@ V-C Comparison with Baselines
To assess FARM under label drift, we compare it against CADE [47] and a controlled ablated variant of our method. CADE and FARM both use an autoencoder-based architecture to learn a latent representation, but they differ in the metric-learning objective: CADE employs contrastive loss, which learns pairwise similarity constraints, whereas FARM uses triplet loss, which explicitly encourages relative separation between anchor, positive, and negative samples. In addition to this external comparison, we include the Percentile variant to isolate the effect of the class-boundary modeling strategy. Percentile uses the same triplet autoencoder and training setup as FARM, but defines class boundaries using class-wise mean centroids and fixed distance thresholds based on the 95th percentile. In contrast, FARM models each class using DBSCAN-derived clusters with cluster-specific dynamic thresholds. The triplet-based variants use the same autoencoder architecture and training settings described in Section IV-B, enabling a controlled comparison of boundary-definition strategies.
@@
Comparative results are presented in Table V. FARM achieves stronger overall performance than both CADE and the ablated Percentile variant, providing a better balance between precision and recall. The comparison with CADE provides an empirical contrast between FARM’s triplet-loss-based latent representation and an autoencoder trained with contrastive loss, while the comparison with Percentile isolates the effect of the boundary modeling strategy, since both methods share the same triplet-based representation. The results suggest that DBSCAN-based clustering with cluster-specific dynamic thresholds improves class boundary modeling and leads to better performance under label drift. The reported values correspond to the mean and standard deviation of precision, recall, and F1 score across experiments on all eight unseen malware families. In each run, one novel family is introduced at a time and evaluated independently.
@@ VI-A Experimental Setup
Once such a cluster is formed, its centroid is treated as a class prototype and incorporated into the classification model. In the case of label drift, where the cluster represents a previously unseen malware family, this prototype is treated as a novel class, allowing the system to classify new instances with only a few examples. In contrast, for covariate drift, where the cluster corresponds to behavioral or structural variations within a previously known malware family, the new prototype is linked to the corresponding existing family label in the training set. In other words, instead of creating a new class, the prototype is added as an additional representative of that known malware family, enriching its characterization in the embedding space. In both scenarios, classification is performed using nearest-prototype matching in the latent space.
@@ VI-C Adaptation under Label Drift
To evaluate the system’s ability to adapt to label drift, we simulate the emergence of a novel malware family that were entirely excluded from the training phase. Samples from each unseen family are introduced one at a time. When a sample is identified as drifted, based on its distance from all known class prototypes, it is stored in a buffer designated for potential novel class discovery. Once the buffer accumulates enough drifted samples, DBSCAN is applied in the latent space to identify cohesive clusters. If a valid cluster is detected, its mean embedding is used to form a new class prototype, which is then incorporated into the classifier. This enables the system to recognize future samples from that family without requiring full model retraining.
@@
VI-C1 Generalization accuracy on unseen classes
To evaluate generalization to previously unseen classes, we conduct standard few-shot classification experiments under multiple $N$-way $K$-shot settings. This evaluation protocol is widely used in the few-shot learning literature to assess whether a model can discriminate among $N$ novel classes given only $K$ labeled examples per class [40, 6, 42].
In our setting, each $N$-way episode is formed by randomly selecting $N$ unseen malware families from the eight families excluded during training. For each family, $K$ labeled samples are used as the support set, and 15 additional samples are used as queries. This protocol reflects a realistic deployment scenario in which the model receives only a limited number of labeled samples from emerging threats and must generalize quickly to classify related variants. Following [40], we run 600 episodes for each $N$-way $K$-shot configuration. The mean classification accuracy and 95% confidence intervals reported in Table VIII are computed over these episodes, providing a statistically grounded estimate of generalization performance across tasks of varying difficulty.
As shown in Table VIII, performance improves consistently as the number of support samples per class increases. The largest gains are observed when moving from 1-shot to 5-shot, while the improvements from 10-shot to 20-shot are more modest, indicating diminishing returns beyond 10 labeled samples per class. In contrast, increasing the number of classes makes the task more difficult, leading to lower accuracy as the model must form finer-grained decision boundaries among a broader set of unseen families.
These results address RQ3 by showing that the learned embedding generalizes well to unseen malware families under limited-label conditions, while also indicating that around 10 support samples already provide a practical tradeoff between adaptation quality and labeling cost.
@@ D. Retraining for Label Drift Integration
While few-shot adaptation enables rapid incorporation of novel malware families, we also include a retraining strategy to support longer-term model integration. This mechanism is applied only for label drift, since the number of samples flagged under covariate drift is typically too small to support reliable model updates. To reduce the risk of overfitting to a very small set of newly observed samples, retraining is triggered only after the drift buffer accumulates at least 100 samples from a novel class. This threshold is used as a practical minimum sample budget for updating the encoder, so that retraining is not initiated from only a handful of potentially unrepresentative instances. In this sense, the threshold reflects a trade-off between adaptation speed and update stability: smaller buffers allow earlier response but may produce unstable latent-space adjustments, whereas larger buffers provide more reliable class representation at the cost of delayed integration. Although the value of 100 is empirically selected, it is intended as a conservative operating point rather than a fixed requirement, and can be adjusted depending on dataset scale, model capacity, and observed drift frequency. Once this threshold is reached, the model is retrained on the union of the original training data and the buffered drift samples, after which DBSCAN is reapplied to rebuild cluster structure and update the corresponding decision thresholds.
Table IX compares the performance of the few-shot adaptation and the retrained model across multiple unseen malware families. The retrained model consistently improves generalization, boosting the average F1 score from 0.85 to 0.94, a relative improvement of 10.6%. This highlights the complementary nature of few-shot adaptation for rapid response and full retraining for deeper integration into the classification system.
TABLE IX: Comparison of few-shot adaptation and full retraining on unseen malware families (label drift scenario).
@@ VII Discussion
FARM shows that combining triplet-based metric learning with unsupervised clustering and few-shot adaptation can provide an effective mechanism for handling both covariate and label drift in malware classification. At the same time, several limitations of the current design should be noted. A central component of the system is DBSCAN clustering in the latent space, which enables the discovery of coherent malware subgroups without requiring predefined family counts. However, clustering performance remains sensitive to the selection of hyperparameters such as $\varepsilon$ and minPts. In this work, $\varepsilon$ was determined using a k-distance elbow heuristic [38], and minPts was set proportional to the latent dimension. While this provides a practical and automated calibration procedure, more adaptive or data-driven parameter selection could further improve drift detection sensitivity, particularly in noisy or structurally complex regions of the latent space.
An illustrative example of this sensitivity is the vobfus family, which exhibited lower drift detection accuracy due to behavioral overlap with training-set families such as fareit and zbot. This overlap caused vobfus samples to be embedded near existing families, resulting in frequent misclassifications. Such cases highlight a limitation of the current embedding space in separating semantically similar but distinct malware behaviors. Representation quality may also be affected by the construction of training triplets. In the current approach, anchor-positive-negative combinations are sampled randomly, without using strategies such as hard positive or hard negative mining. Although this simplifies training and avoids overemphasizing extreme cases, it may limit the model’s ability to refine decision boundaries in ambiguous or tightly clustered regions.
A further limitation concerns the evaluation protocol. FARM is assessed here in a drift-response setting, where the model is first trained on an initial family distribution and then updated through prototype insertion and later retraining as drifted samples accumulate. While this setup reflects an operational malware adaptation scenario and includes a longer-term consolidation stage, it does not constitute a full continual learning benchmark with many sequential tasks, replay strategies, or explicit forgetting measurements. Therefore, the present results should be interpreted as evidence of effective drift-aware adaptation rather than as a claim of superiority over established continual learning methods.
Another scope limitation is that the present study is instantiated on EMBER feature representations for Windows PE malware. Although the overall framework is compatible with other feature extractors, the quality of the learned latent space and the resulting drift boundaries will depend on the representational properties of the chosen features. Extending the framework to alternative static, dynamic, or hybrid malware representations remains an important direction for future work.
Overall, the results suggest that FARM provides a practical basis for drift-aware malware adaptation, while also revealing several areas where the current formulation can be strengthened. In particular, improving clustering calibration, refining triplet construction, and evaluating the framework under broader sequential adaptation settings would further clarify its robustness and generality.
@@ VIII Conclusion
This work addresses the challenge of maintaining malware family classification performance under continuous concept drift. By integrating metric learning, few-shot adaptation, and conditional retraining, FARM provides a unified framework for identifying and responding to both evolved and novel malware families. The use of a triplet autoencoder enables drift detection without relying on softmax confidence scores, while few-shot learning supports rapid adaptation using only a small number of labeled samples. In addition, the inclusion of a buffer-triggered retraining mechanism allows FARM to incorporate newly observed data into the model and refresh its latent representation for longer-term updating. Experimental results show that FARM improves classification performance across both covariate drift and label drift scenarios, indicating its effectiveness in temporally evolving settings. These findings suggest that FARM can serve as a practical approach for drift-aware malware family classification in dynamic environments. Future work may extend the framework to handle packed malware, which introduces additional drift through obfuscation and poses challenges for static analysis. Another important direction is to improve the interpretability of drift decisions by identifying the underlying changes that trigger detection, thereby increasing transparency and supporting more informed operational responses.
Code Availability: The source code and scripts developed for this paper are available at https://github.com/numanhg/farm.